Calculate the molarity of a solution that has 17 grams of sodium chloride dissolved in 250 ml of water.


1
Expert's answer
2021-04-15T03:22:05-0400

m(NaCl) = m/M = 17g/58.5g/mol = 0.29 mol

Molarity = n/V = 0.29mol/0.25L = 1.16 mol/L = 1.16 M